OpenStudy (anonymous):

find the number of all possible triplets(a1,a2,a3) such thata1+a2sin2x+a3sin^2x=0.

OpenStudy (anonymous):

is this ur question?\[a_1+a_2 \sin 2x+a_3 \ \sin^2 x=0\]

OpenStudy (anonymous):

\[\frac{ -2a1 - a3 }{ 2 } = a2\sin2x - \frac{ a3 }{ 2 }\cos2x \]

OpenStudy (anonymous):

Clearly only visible condition on a1 , a2, a3 is that: \[-1 \le \frac{ 2a1 + a3 }{ \sqrt{4(a2)^2 + (a3)^2} }\]

OpenStudy (anonymous):

\[\le1\]

OpenStudy (anonymous):

MAybe I'm missing something. As this still gives infinite cases.

OpenStudy (anonymous):

i think its supposed to be \[a_1+a_2 \cos 2x+a_3 \ \sin^2 x=0\]more likely

OpenStudy (anonymous):

Okay. That is probable.

OpenStudy (anonymous):

just a thought :)